Sol Summary – November 20th

Sol Summary Report

November 20, 2017

…and new civilizations…

Shannon Rupert, MDRS Director

Mission Status: Still not in sim.  Still just Brandon and I although
we had many visitors today working on the preparations for the arrival
of the RAMM, our new repair and maintenance module, which lands on
station at 10 am tomorrow.

Sol Activity Summary:  I spent some time in the GreenHab this morning,
and this afternoon Brandon and I parted ways: he went to explore Coal
Mine Wash and then got water and I went to the canyon I discovered
last winter and explored.  I found evidence of a drying inland sea and
Brandon found the pond was already frozen in Coal Mine Wash.  He also
found the overlook on Skyline Rim.

Look Ahead Plan:  RAMM arrives, we clean and prep for Crew 183.  I
just realized that Crew 183 won’t be here tomorrow but rather the next
day. But I also just today realized that this week is Thanksgiving in
the US, so my timeline is all messed up!

Anomalies in work:  Other than not being in sim, it was just another
day on Mars.

Weather:  Calm, sunny and cold…again.

Crew Physical Status:  We are both fine.

EVA: Brandon out of sim in the HabCar to Coal Mine Wash, me in rover
to Discovery Canyon (for lack of a better name).

Reports to be file: Sol Summary and Operations Report.  No photos.

Support Requested: None at this time.

Operations Report – November 20th

Crew 182+ Operations Report  20NOV2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: same as yesterday, except the generator is
running well again

Notes on non-nominal systems:  Generator run all night last night, so
we don’t know if the last failure was due to the diesel or needing
oil.

Generator (hours run): 24

Solar— SOC 100 %

Diesel –  63 %

Propane –  not known psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –   0 gallons.

Water (trailer) –  0 gallons.

Water (static) –  150 gallons (Brandon is currently bringing 100
gallons but we will need to make a couple more trips tomorrow.

Trailer to Static Pump used –  no

Water (loft) – Static to Loft Pump used –  not known

Water Meter:  not checked

Toilet tank emptied: not that I know of

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) All 5 were moved to make
space for the big equipment preparing for the RAMM’s arrival tomorrow

Oil Added? no

ATV Fuel Used:  0Gals

# Hours the ATVs were Used today: 5 min each

Notes on ATVs: The 300’s battery is low, will need to jump it tomorrow

Deimos rover used:  yes

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Sojourner rover used:  ASSIGNED TO DIRECTOR, not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used:  Not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Opportunity rover used:  not yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

Curiosity rover used:  no yet ready

Hours:

Beginning charge:

Ending charge:

Currently charging: no

HabCar used? Yes

General notes and comments:  The new rovers need to be put together
and the seats modified (except for Sojourner).  Not yet started on
this project.  Will charge the other three tomorrow.

Summary of internet:  Internet was upgraded at Shannon’s, but the
phone does not work properly.  Internet seems as good as before it
went bad, but not blindingly faster even with the upgrade.  Wifi is
turned off to this service.  Hab internet is good.

Summary of suits:  n/a, not used

Summary of Hab operations: The forced air furnace is still
periodically blowing cold air.  Need to clean the filter tomorrow then
order some as soon as possible to see if that is the problem.

Summary of GreenHab operations:  I moved the beans to the wooden bench
and provided cages for their continued growth.  Other plants were
organized.  I prepped 10 pots for moringa seeds.  Ordered further
supplies: timer for lights, environmental monitor, herb and lettuce
seeds with planters

Summary of ScienceDome operations: Ordered two floor lamps for evening
light.  We will have to clean the lab tomorrow as it is not in
changeover condition.

Summary of RAMM operations:  The area was leveled and the swale was
dug so that water will go around the building during rain and runoff.
The RAMM left California this evening and will be onsite at 10 am.  It
will be lowered onto a railroad tie frame to keep it off the ground.

Summary of health and safety issues:  Nothing to report

Questions, concerns and requests to Mission Support:  Nothing, thank you.

Sol Summary – November 19th

Sol Summary Report

November 19, 2017

Seek out new worlds…

Shannon Rupert, MDRS Director

Mission Status: We are not in sim. Brandon and I are the only two
here.  Les and his team departed early.  We are awaiting Crew 183 on
Tuesday, when they will begin their one-week mission.

Sol Activity Summary:  We did some map/road checking today, exploring
road 1574 and Cactus Road.  Updates will be in my soon to be released
guide to MDRS roads.

Look Ahead Plan: Tomorrow HudgesNet comes to fix my service and we
begin work on the pad for the new RAMM, which will be installed on
campus on Tuesday.  I will also order winter blend diesel, because I
believe it is now cold enough that the summer mix is what shut down
the generator last night.  We also need to get water, ethanol free gas
and work on the new rovers.

Anomalies in work:  Not a sim day.  Generator did not run all night,
but has run all day.  I suspect the summer blend diesel is the issue.
Tried to get lower deck heater to work this morning but it did not.
Will continue troubleshooting.  Water was frozen in the am but thawed
by noon.  We need to get the tank and pipes insulated soon.  I will
contact the contractor.

Weather:  Calm, sunny and cold.

Crew Physical Status:  We are both fine.

EVA: Out of sim EVA to check out northern routes.

Reports to be file: Sol Summary and Operations Report, which was sent
earlier today.  No photos will be sent.

Support Requested: NOTICE:  Effectively immediately and until I say
differently in the spring, the heater at the Hab will stay on as
follows:

Temperature 72 degrees

Fan on auto

Heater on ON

The controller should not be adjusted at all by any crew.  When you
adjust it, we lose the heater and it just blows cold air, but if you
just leave it alone, it works perfectly.  If the crew is cold, they
should turn on the secondary heater (once I get it running!)

Please find out what type of filter the heater uses and let me know so
I can order some.  Thank you.

Operations Report – November 19th

Crew 182+ Operations Report 19NOV2017

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: water heater, lower deck electrical outlets, one
suit charger, generator, Hab water system, front airlock window, top
hatch

Notes: Water heater needs to be replaced. It was supposed to be
replaced two Mondays ago, but when the contractor came to do it, the
crew did not open the door

Lower deck electrical outlets: Electrician is being hired to look at this.

Suit charger: Broken by crewmember. An additional three suit chargers
were not working when the suits were delivered. Contacting NorCal
regarding replacement.

Generator failed in the night. Believed to be low oil. Will monitor.

Last night when I went into the Hab, the water pump was running
non-stop and the system was empty. After I refilled the tank to ½ and
troubleshooted, I determined that someone just ran the system dry. I
reprimed the lines and the pump began functioning normally.

The front airlock windown came out this summer, and during the refit
and again the night before last. Need to find a way to secure it.

The new top hatch is not secure in high winds. Need to find a way to secure it.

Generator— Turned off yesterday am at 8 am, turned on last night at 6
pm, failed during the night, was restarted after adding oil at 7:45
am. SOC was at 61% so I’m going to let it run all day. We need to
replace the summer blend with a winter blend and we did not use the
amount of fuel we thought we would because the generator was out of
service all summer. Note: only use the generator specific oil.

Solar— SOC was at 61% this morning. Will monitor throughout the day.

Diesel – 65%

Propane – n/a psi. Amerigas was contacted and propane was ordered.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 0 gallons. Note
this should be rewritten on the template as Ethanol Free gas. This is
the only gas that should go in the ATVs. It is only available at
Stan’s Chevron. In addition, the only oil that should go in the ATVs
is an ATV specialty oil.

Water (trailer) – 0 gallons.

Water (static) – ~ 200 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: no (at least not by me)

ATV’s Used: All ATVs except the 300 were used today. All 5 have been
serviced and repaired for the season. The 300 still needs its parking
brake repaired, parts are on order. The 350s all have new seats so
none of the machines has a ripped seat. We will be using one today.

Oil Checked: no

Atv # Fuel Used Gals: Not none. We will go get gas toady.

Tires Status: Fine

# Hours the ATVs were Used today:

Deimos used: Yes

Hours: nor reported

Beginning charge:

Ending charge:

Currently charging: yes

HabCar used? Yes. We got one load of water in it today.

Notes and Comments: The four new rovers are charged but still need to
be put together and their seats modified. Progress will be reports as
I get to them.

We are using so much less water with the new toilet.

Summary of internet: Also needs to be added to template. Not sure
about Hab’s internet.

Summary of suits: This needs to be added to the template. All suits
are operational, 6 suits in EVA room. Two exploration suits in EVA
room, only one charger. Prototype exploration suit in EVA room. Two
exploration suits in Science Dome with no working chargers. I found a
bottle of soda lime on the stirrer in the lab and it was on. I turned
it off (hope it wasn’t Brandon’s, just thought about this).

Summary of Hab operations: Noted several failed systems and they are
noted above. The cement pad out the rear airlock needs some serious
cleaning.

Summary of GreenHab operations: Heating at 50 degrees F. It’s lovely
in there. Plants are looking good. Hose has a leak and freezes at
night. Needs to be modified to a very short hose that is taken in
every night or water taken directly from freeze free faucet. Crews
need to be sure they push the level all the way down and use it daily
to prevent freezing. Caution should be taken to not spill any water
on the ground as it will turn to dangerous ice,

Summary of ScienceDome operations: The lab is pretty much a mess.
Brandon is still doing his work, but lab will require cleaning before
next crew. Shop vac was not used nor put together.

Summary of health and safety issues: Still need alarms, will check to
be sure they were ordered.

Questions, concerns and requests to Mission Support: Please update
this template with the additions indicated. Thank you.

Crew 182 – End of Mission Summary

Team Peru V:

Commander/ Green Hab/ Health and Safety Officer: Atila Meszaros (Peru)

Executive Officer/ Crew Journalist/ Scientist: Camila Castillo (Peru)

Engineer Officer: Carmen Atauconcha (Peru)

EVA Officer/ Crew Geologist: Brandon Fergurson (USA)

Crew member: Julio Rezende (Brazil)

Crew 182

 

The Mars Society Peru Chapter sent Team Peru V (Crew 182), conformed by a multidisciplinary group. Their rotation was scheduled for November 4th (the day Carmen ate those burgers without us) – 18th 2017.  The main goal of the crew was to develop research in their different fields at the MDRS, achieving their specific goals. The multidisciplinary approach of the crew proved to be valuable during the mission.

 

During the mission, the following research activities took place at MDRS:

 

  1. Effect of Streptomyces sp. Isolated from mineral cultures on radish plant development in analog martian soil: Soil was collected around the MDRS location to use it for radish crops. The strain used at the inoculation was isolated from mineral cultures, which are also an extreme environment. The main objective is to prove the effect of this strain in crops in martian analog soil. The main goal of this research was unachieved, but soil samples will be taken to Lima (Peru) for further experiments.

 

  1. Resistance of Peruvian Altiplano’s crops to martian analog soil: Soils with different compositions where collected on the surroundings of MDRS and on the Salt Wash Member of the Morrison Formation in order to prove the resistance of Peruvian crops and mustard (as control) to mars analog soil. The main goal of the project wasn’t achieved, mostly because an incident during #7 EVA. However, the research will continue on Lima (Peru) using the martian analog soil and two more altiplano’s seeds.

 

  1. Incidence of consumption of kiwicha cookies in the loss of muscle mass that people living in the analog of Mars experiment: I prepared cookies of kiwicha on Peru, kiwicha is an andin grain that has enormous amounts of protein. Because of this characteristic of the kiwicha grain, my cookies have 10% of protein per portion. During the time that I spend in the rotation, I had to take notes of the mass muscle index. So, I gave the cookies to half of the crew, two units per day. Also, every 4 days I took notes of their weight. With this data, I am going to compare the data of the crew member that ate the cookies and the ones that do not ate the kiwicha cookies.

 

  1. Properties and Composition of Mars Analog Regolith at MDRS: Regolith samples were collected from different areas within the MDRS area. The study focuses specifically on the Morrison geologic formation. The majority of the samples are from the brushy basin member of this formation. The goal of the project is to classify the soil properties including: soil texture, classification, and composition. The project will continue during the next week.

 

  1. Sustainability in Mars research stations and extraplanetary settlements: This research searches to answer the question: The Mars Desert Research Station (MDRS) operation can be more sustainable? It is evaluated how environmental, economic, social and personal sustainability issues are presented in the research station and how the MDRS activities would collaborate to Sustainable Development Goals (SDGs), proposing some guidelines to sustainability. It is also important to ask: the results related to Mars would be applied to build a self-sustainable habitat in Earth, mainly in areas affected by climate change, as deserts and semiarid regions as can be seen in Brazilian Northeast (Habitat Marte)? Reviewing the previous research done at MDRS not was identified any research related to sustainability. Because of that, this research presented a high impact to MDRS and Mars research. It is a challenger identifies the main dimensions that would be considered to evaluates a Mars research station in terms of sustainability: this is the great relevance of this research for the future design of Mars settlements.

 

Sol Summary – November 17th

MDRS Daily Summary Report for Sol 10

Summary Title: Is this the real life?…

Mission Status: No incidents to report today, except that our crew biologist was kind of trippy today in the EVA session. She had to many pancakes.

Sol Activity Summary: Or is this just fantasy? Today we met with Survivorman, also known as Les Stroud. A pretty calm guy with really interesting experiences around the world. Both he and the cameramen were great and we had a great time. But first, let’s explain how our day started: It was just a normal day… expecting the arrival of Survivorman, but a normal day none-the less. We asked again about the pancakes and all those things that became inside jokes around the crew. And then we realized: Man, this will be our last night in MDRS, right? The thought is kind of frightening… specially for me, because I need to do an exposition in my University about some paper I haven’t even read. I would like to stay in MDRS at least a little bit longer. The bacterias I cultivated here have so many different shapes and colors, it is quite exciting for me as a wanna-be microbiologist.

After the philosophical moment, Survivorman arrived! Fortunately, we planned an EVA today, one where yours truly participated. We went to the Lith canyon. The new ATVs were great, but the trekking there got tiring after a while. We were so tired that some crewmembers here made some associations with Igor the donkey of Winnie the Pooh (Get it? There’s a place called after Pooh, so we think about Winnie the Pooh characters… ok, maybe that’s another inside joke). The cameramen started filming us, and I must say they got nice shots. That Lith Canyon place is absolutely beautiful and so interesting. It had so much geological material and the plants there were thrilling! I will always be amazed by how life still goes on even in such places as deserts.

We arrived to the hab with the filmcrew and decided to take some lunch. But, before that, we decided to do what we were expecting since day one: Eating the russian space food that they sent so gently to us. We were waiting for Survivorman to make this even more epic. And there we were, holding our toothpaste-like tubes to our mouths, tasting the meat and the apple pure. I must say that I liked both, but some of the guys just couldn’t get around with the new flavour. Anyways, it was quite a shocking experience.

After a nice lunch at the hab, we finally got some rest. I had to do some work at the lab, but the guys stayed here to help bring in the new food for the next crew.

I still can’t believe this will be our last day at MDRS. But I’m glad I had the chance to get to know the crewmembers, as well as Survivorman, the people of Mars Society we got to know (yes, even the Capcoms) and, of course, the desert.

Maybe a proper thank you wouldn’t be enough, y’know? But at least I have a story to tell my grandchildren in the future, with some good pancakes.

Look Ahead Plan: Our homes are waiting in our respective countries, so as long as we arrive on time and safely, that’s the only plan we have.

Anomalies in work: Nope.

Weather: Windy and with some rain.

Crew Physical Status: Very well.

EVA: Most of our EVA is explained in the summary. We collected some samples of geological interest. The cameramen did some shots of our activities during this EVA.

Reports to be file:

Commander Report
EVA Report
Science Report
Engineering Report
Green Hab Report
Final Mission Report

Support Requested: Nope.

Camila Castillo, Crew Biologist and Journalist

GreenHab Report – November 17th

Sol 14 – 17/11/201

Green Hab Officer Atila Meszaros

Ambient: heating 10°C

Ambient with window/door open: door opened at 09:00; closed at 14:35

Shade cloth on/off: on

Average temperature:

Time 08:05 14:35 18:16
Temperature inside (°C) 29 15 16
Temperature outside (°C) 16 8 6
% Humidity 16 16 16

 

Changes to crops:  Bigger than ever!

Daily water usage for crops: 2 galons

Time(s) of watering for crops:  08:05 and 18:00

Changes to research plants: controls keep growing!

Daily watering and amount of water used: 1 ½ L approximately

Narrative:

Hi, Veronica, Atila in here, again. I hope you are having an amazing night.

Today is our last simulation day! Jeez, times really fly. I’m going to miss those little guys (I’m talking about the crops, of course, the crew too, tough). They are just huge.

We already ended simulation, and this is the final Green Hab Report, however tomorrow I’ll water the crops twice at 08:00 and 13:00.

I couldn’t obtain any conclusive results from my project probably because of the incident of #7 EVA (I wasn’t able to water the research plants that Sol). However, I’m aiming to continue my research with the analog soil on my university. I hope I will be able to let you know how that worked.

I forgot to mention in my last Green Hab Report, that, with help of the whole crew, the Green Hab is shining! We put our bodies and souls on that.

I think that’s pretty much all, if you need more information about the GH I will be glad to send it to you.

Thank you for being there!

Have a beautiful night on Earth.

Support/supplies needed: none

Operations Report – November 17th

Crew 182 Operations Report  17Nov2017

SOL: 14

Name of person filing report:  Carmen Atauconcha

Watching the beautiful hills on Mars, I truly could say that it has become my home. Mars with its mysterious shadows and shining soils has conquered my hearth.

This is my final engineering report:

Non-nominal systems:

WATER HEATER: The new water heater is not installed, yet. So, good luck to the next crew¡

CHARGERS OF THE SUITS: Two of the chargers are working, but we moved them and now it is hard for us to connect them, again.

THE DOOR AIRLOCK’S WINDOW: This day was very windy. I think Mars does not want to leave us. The fact is, because of the force of the wind the window went out. Now, it is secure inside the Hab, next to the astronaut costumes.

COMMUNICATION RADIOS: Three radios are not working. The first one because it is broke. The other two are not transmitting signal.

 

Generator—  It worked the whole night since 16:50pm of 16Nov2017 until 9:20 am of 17Nov2017.  Turn on at 17:00 pm 17Nov2017.  Operand.

Solar—

READING FROM 17 NOV 2017

 

TIME SOC READINGS
9:20 am 100%
10:20 am 100%
11:20 am 100%
12:20 pm 99%
13:20 pm 95%
14:20 pm 91%
15:00 pm 89%
15:20 pm 85%
16:00 pm 83%
16:20 pm 81%
17:00 pm 76%

 

Diesel – 50 %

Propane – 49 psi

Gasoline (5 Gallon containers for ATV) –  2 ½ gallons.

Water (trailer) –  0 gallons

Water (static) – 300  gallons

Trailer to Static Pump used –   No

Water (loft) – Static to Loft Pump used –  yes

Water Meter:  60 gallons

Toilet tank emptied:  yes

ATV’s Used: Yes

Oil Checked:  Yes

Atv # Fuel Used Gals:  4  gallons

Tires Status: Perfect

# Hours the ATVs were Used today: 3 hours

Deimos used:  Yes

Hours: 3

Beginning charge:  – 15:40pm

Ending charge:  –

Currently charging: Yes
HabCar used? Yes

Notes and Comments:

There are no words to express our experience on Mars. Each day has formed our relationship between each crew as a family.  About Mars, what can I say? I am in love of it.

 

Summary of Hab Operations:
In the morning we covered the hole that is in the highest part of the roof. Thus, the wind moved the hollow coverage and it is open again. We will fix it tonight again.  The toilet is starting to smell better than it was in our firs days on Mars.

The Hab is cleaned and organized. Be careful with the secondary airlock because it is wonky.  The heater is working very well.  The area of the tools is organized, please keep it like that.

Summary of GreenHab operations:

The tools and equipments in the greenhouse are organized and cleaned.  Take care of the plants because they are growing.  The greenhouse is like the lung of the station.

Summary of ScienceDome operations:

It is necessary to install a lamp there. The autoclave and the rest of equipments are working very nicely. Be careful while you use the autoclave, please read the instructions in of the manual that is in that laboratory. Without science we won’t be on Mars.

Summary of health and safety issues:

It is hard to bring the food from the first floor of the Hab to the second floor. This fact is because the stairs are very steep. Thanks for the new stairs. Health and safety are first.

This experience has ended and my report, too. How the water moves on Earth, the time flies on Mars. We only have the hope to understand how the universe works. You will be always in my heart Mars.

Science Report – November 17th

Science Report
Camila Castillo
17 Novermber 2017

The research project about the effect of Streptomyces spp. strain could not be finished properly. The first problem was related with the lack of medium needed for finishing this project. Nutrient broth and Nutrient agar were the ones I worked with, so there are better results with these. I decided to buy these elements via Internet, but I had to wait a week. After that week, there was little time to start the project. I decided to start working right away, but sadly, the medium ended up contaminated. I will need to check if the transportation of the strain was done properly, though. It would seem that that was the main problem. Maybe it couldn’t resist as much as I expected. Microbiology projects tend to have these kind of problems, though. As a beginner in the microbiological world, maybe I still need to learn some more about microbiology techniques to work an entire project in just two weeks! I will be taking with me some soil samples to Peru. I hope that the enviroment provided by my lab at Lima will help me to finish this project.
Despite the kind of failure of my main project, I managed to learn some things about martian agriculture. First of all, the soil isn’t as simple as I expected. It hardens and it prevents the water to filter through it. It has little to non-organic matter. Somehow I thought that my mere strain could help with these problems, but I realized that this is a problem I will need to adress if I want to finish this (which I will). I tried to search for a technique that didn’t involved mixing two types of soil, but both didn’t work. It seems that in the end I will have to do the mixing.
However, I also managed to do some projects besides my main one. I tried to see the effects of the dishwater on the seeds we had, and I managed to do a nice diagram about some samples and their growth in a couple of days. I also decided to describe the microbes in different plant soils we managed to collect on the EVAs. I’ve been amazed by the shape and forms I managed to get on my plates. I will try to take some of these samples to Lima and process them, to know exactly which microbes are in there.

EVA Report – November 17th

EVA #10 Report

Today’s EVA began at 10:00 AM, involving Camila, Julio, and I. The Survivalman crew was filming our EVA, including the planning process, communications tests, putting on the spacesuits, and our decompression. The goal of our EVA was to go to Lith Canyon, collect three regolith samples, and return to the Hab. Each of the EVA crew boarded an ATV, with Les Stroud on his own ATV, and his camera crew on Deimos. I had Camila lead the way out so we could keep our ATV caravan nice and tight. The trip there took about an hour, because they needed footage of us along several different stretches of Cow Dung Rd. We would hold our position as they moved ahead, then we would move forward once they signaled that they were ready.

Once we arrived at the end of the Lith Canyon road, we proceeded on foot for about a kilometer westward. We had to be careful with our route finding as this was an area that none of us had been to before. We made sure to photograph the landscape at certain points so it would be easier to find our location later. Once we descended into the canyon, we made our way to an exposure with red and white regolith layers, taking one sample from each type. Once finished we walked further down the canyon where there were several eolian rock formations. There we collected our last regolith sample from a white exposure, turned around, and headed back the way we came. We spent about 2 hours walking in the Lith Canyon area, returning to the ATVs by around 1:20 PM. They didn’t need any more footage for the trip back, so we made our way back to the Hab by around 1:50 PM.

Today’s EVA was abnormally warm, especially all the walking we had to do. This was the longest EVA that I’ve had to do on foot, and was very challenging. I’d estimate that we walked a total of about 2 km over uneven terrain. It’s amazing how heavy and uncomfortable the suits felt on the first EVA when we just walked around the Hab. Now I feel a lot better in the EVA suit, even when covering much longer distances. Knowing this, I’m glad we didn’t plan any really long EVAs during our first few days of Sim, as it does take some time to get used to those suits. I’d definitely recommend that crew members take their time getting used to the suits, as their bulkiness can be surprising.

For our EVA, we managed to obtain the regolith samples we needed and the Survivorman crew was able to record our EVA operations.

This is all I have to report for today’s EVA.

Have a great night Mission Support!

 

Prepared by Brandon Ferguson

EVA Officer